OpenStudy (anonymous):

can someone help me please... 7/8 divided by 12

OpenStudy (anonymous):

Hi bnhant i know how to do this.

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

I'll draw.

OpenStudy (anonymous):

no prob.

OpenStudy (anonymous):

\[\frac{7}{8} \div12 = \frac{7}{8} \div \frac{12}{1} = \frac{7}{8} \times \frac{1}{12} \]

OpenStudy (anonymous):

\[= \frac{7}{8 \times 12} \]

OpenStudy (anonymous):

\[= \frac{7}{96} \]

OpenStudy (anonymous):

is that the final answer 7 over 96

OpenStudy (anonymous):

yes like eigen showed you

OpenStudy (anonymous):

yes. do you understand the steps?

OpenStudy (anonymous):

yes i am older and forget how to do all of this and im checking my daughters work for her that she did

OpenStudy (anonymous):

Eigen, when you can, could you please go back to my question? Not to be annoying or anything, but I need help badly.

OpenStudy (anonymous):

sure :D